OpenStudy (anonymous):

Domain question? f(x)=4/x^2

OpenStudy (anonymous):

I have choices. Hang on. >:P

OpenStudy (australopithecus):

What number can you not sub in as x?

OpenStudy (australopithecus):

What number will give you a undefined value for f(x)

OpenStudy (anonymous):

I guess zero isn't an option. Choices: A) x ≥ 0 B) All real numbers except 0 C) All real numbers except 3 D) All real numbers

OpenStudy (australopithecus):

Any guesses at all?

OpenStudy (anonymous):

It's B, right?

OpenStudy (australopithecus):

Yup

OpenStudy (anonymous):

f(x) = (x - 2)^2 + 2 What about this one? It's range instead of domain... I just can't seem to get range or domain. :c @Australopithecus

OpenStudy (anonymous):

Yay! C:

OpenStudy (australopithecus):

Domains are simple, Just remember the following math rules -You cannot take the logarithm of a number equal to or less than zero, For example: log(-1) =/= undefined (if you are curious as to why this is just remember log(x) = y is the same as 10^y = x try to make x negative in this form, you cant. -You cannot divide by zero so, For example 0/0 = undefined 1/0 = undefined 0/1 = 0 -You cannot take the square root of a negative number, you can take the square root of zero For example, (-2)^(1/2) = an imaginary number (-3)^(1/6) = an imaginary number (0)^(1/66) = 0 Another hint always check the exponent on the variable x, if it is even remember all numbers subbed into it will give you a positive, for example, (-2)^2 = 4 (-2)^3 = -8 (-2)^4 = 16 Im pretty sure this covers domains

OpenStudy (anonymous):

Oh, gosh. Thanks. C: Is it pretty much the same thing for range?

OpenStudy (australopithecus):

Well not really, for the range I normally check the extreme values of x in the domain, or I graph the function, or if it is a polynomial you can find its vertex (Look up guides on youtube)

OpenStudy (anonymous):

Awesome. Thanks. c: God bless you!